OpenStudy (anonymous):

Someone explain a math lesson to me. My teacher named the assignment "Factor with X," but the assignment doesn't have to do with factoring alone. It also involves brackets and grouping, I think he mentioned? So far, I think the lesson is called "Factoring trinomials with a common factor." But I'm still a little unsure.

OpenStudy (anonymous):

you got a specific problem to solve?

OpenStudy (anonymous):

7b^2 + 46b + 24

OpenStudy (amistre64):

if its rational then given: ax^2 + bx + c we can start factoring it as ac = k, such that k has factors that add to b

OpenStudy (amistre64):

spose ac = k = mn, such that m+n=b then we can factor this as: (x+m/a) (x+n/a) and we are left with 2 steps to clean it up

OpenStudy (amistre64):

7(24) = 168 1,168 2,84 4,42 ... 4+42 = 46 sooo (x+4/7) (x+42/7) simplify, if theres still a fraction left, put the bottom in front (x+4/7) (x+6) (7x+4) (x+6)

jimthompson5910 (jim_thompson5910):

you can use the quadratic formula to solve 7x^2 + 46x + 24 = 0 for x to get these two solutions x = -4/7 or x = -6 then transform each equation to make 0 be on the right side 7x = -4 or x = -6 7x+4 = 0 or x+6 = 0 Finally, you can use the zero product property to get (7x+4)(x+6) = 0 This shows us that 7x^2 + 46x + 24 factors to (7x+4)(x+6)

OpenStudy (anonymous):

I don't understand where you got -4/7 and -6 from. And I'm not familiar with the quadratic formula.

jimthompson5910 (jim_thompson5910):

if you haven't learned about the quadratic formula yet, then don't worry about it

jimthompson5910 (jim_thompson5910):

it's something you'll learn later on in your algebra course it's a formula to solve things like 7x^2 + 46x + 24 = 0 for x

jimthompson5910 (jim_thompson5910):

here is a page with a few examples http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

So, I would substitute x for 0?

jimthompson5910 (jim_thompson5910):

no you would plug a = 7, b = 46 and c = 24 into the quadratic formula

jimthompson5910 (jim_thompson5910):

then simplify

OpenStudy (anonymous):

Why is the plus sign underlined? What does it mean?

jimthompson5910 (jim_thompson5910):

it represents "plus or minus" when we say something like "5 plus or minus 3" we mean 5+3 or 5-3

jimthompson5910 (jim_thompson5910):

so what's really going on is that we have two equations

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):